Ball-Bearings, Inc. produces ball bearings automatically on a Kronar BBX machine. For one of the ball bearings, the mean diameter is set at 20.00 mm (millimetres). The standard deviation of the production over a long period of time was computed to be 0.150 mm. What percent of the ball bearings will have diameters 20.27 mm or more?
Answer #1

Solution :

Given ,

mean = = 20.00

standard deviation = = 0.150

P(x > 20.27) = 1 - P(x<20.27 )

= 1 - P[(x -) / < (20.27 -20.00) / 0.150]

= 1 - P(z <1.8 )

Using z table

= 1 - 0.9641

= 0.0359

answer =3.59%